If NM=8x-14 and JK=x^2+1 Find JK, I know that since the shape is a rectangle you'd start out with 8x-14=x^2+1 but how do I sort out the x^2?
Answer by venugopalramana(3286)

If NM=8x-14 and JK=x^2+1..WHAT ARE THESE EXPRESSIONS GIVEN FOR JK AND MN ? ARE THEY EQNS.OF JK AND MN?IF SO JK IS NOT A STRAIGHT LINE.
Find JK,..WHAT IS MEANT BY THIS ?.JK IS ALREADY GIVEN ABOVE!
I know that since the shape is a rectangle..HOW SINCE JK IS NOT A STRAIGHT LINE JKMN CANNOT BE A RECTANGLE
you'd start out with 8x-14=x^2+1 but how do I sort out the x^2?
SO PLEASE DESCRIBE PROPERLY THE PROBLEM AND WHAT IS TO BE FOUND?GIVE A SKETCH OTHERWISE...